With the current, a motorboat travels 60 mi in 3 h. Against the current, it takes 2 h longer to travel 60 mi. Find the rate of the current.
Answer by Paul(988)

Let y be the speed of the boat
Let x be the speed of the current
2 hours longer = 5 hours total
FIRST EQUATION:
3x+3y=60
y=20-x (Subsitution)

SECOND EQUATION:
5x-5y=60
y=12+x
SUBSITUTE
20-x=12+x
2x=8
x=4
y=20-4
y=16
Hence, the speed of the boat is 16mph and the speed of the current 4mph.
Paul.